# Learning How to Dynamically Decouple

### v1

@@ Abstract
Current quantum computers suffer from noise that stems from interactions between the quantum system that constitutes the quantum device and its environment. These interactions can be suppressed through dynamical decoupling to reduce computational errors. However, the performance of dynamical decoupling depends on the type of the system-environment interactions that are present, which often lack an accurate model in quantum devices. We show that the performance of dynamical decoupling can be improved by optimizing its rotational gates to tailor them to the quantum hardware. We find that compared to canonical decoupling sequences, such as CPMG and XY4, the optimized dynamical decoupling sequences yield the best performance in suppressing noise in superconducting qubits. Our work thus enhances existing error suppression methods which helps increase circuit depth and result quality on noisy hardware.
@@ I Introduction
Figure 1: Illustration of LDD. The left and the right block indicates a quantum device and a classical computer, respectively, working in tandem. A DD sequence parameterized by $\vec{x}$ is added to idling qubits in a given quantum circuit, here exemplified by a SWAP gate on $(q_{1},q_{2})$ followed by a CNOT on $(q_{0},q_{1})$. The arbitrary single-qubit rotations $R_{\vec{x}}$ in the DD sequence are realized by three parameterized virtual $Z$-gates (circular arrows) and two $\sqrt{X}$ gates (pulse with in-phase and quadrature shown in light and dark blue, respectively). The noisy output state $\rho(\vec{x})$ is compared to the expected ideal state $\rho_{\text{ideal}}$ through the cost function $J(\vec{x})$. A classical optimization algorithm minimizes $J(\vec{x})$ to find the parameters of the gates in the DD sequence that best fit the circuit.
@@
In recent years, DD has become a major error suppression method for noisy superconducting quantum computers [16]. Indeed, the method is easy to apply as a simple transpilation pass that inserts delays and pulses into a quantum circuit. Furthermore, simple sequences such as $X-X$, where $X$ is a $\pi$-rotation around the qubit’s $x$-axis, already yield excellent results [20]. More elaborate sequences, such as staggered $X-X$ [21] and staggered $XY4$, improve, for instance, the execution of dynamic circuits by cancelling cross-talk [22]. Crucially, the performance of a DD sequence depends on the interactions present in the quantum hardware. In superconducting qubits [23], a good model of these interactions is typically not known, a case familiar to optimal control [24, 25, 26, 27] that can be overcome with closed-loop optimization [28, 29]. Similarly, it is possible to tailor DD sequences to the quantum hardware at hand by learning them through genetic algorithms [30]. However, in hardware, DD pulses are not perfect, potentially introducing additional noise and errors through the controls that implement the DD pulses, thereby diminishing the quality of the designed DD sequence.
@@
To overcome these limitations, in this work we tailor the DD sequences to the hardware and quantum circuits to execute. This is achieved by optimizing the rotational angles of the gates in the DD sequence in a closed-loop with the quantum hardware. A classical optimizer is fed the cost function value that is reconstructed from quantum samples and that is sensitive to the quality of the DD pulses, see Fig. 1.
@@
This manuscript is structured as follows. We introduce in Sec. II two commonly employed DD sequences CPMG and XY4. Next, in Sec. III we develop the theoretical framework of how optimal parameters in DD sequences are found on quantum hardware, which we refer to as learning dynamical decoupling (LDD). We demonstrate in Sec. III.2 the utility of LDD on IBM Quantum hardware by comparing the performance of LDD to CPMG and XY4 to suppress noise in two experiments. We show that LDD outperforms CPMG and XY4 for suppressing noise present during mid-circuit measurements and noise resulting from increasing the depth of a quantum circuit. We conclude in Sec. IV.
@@ II Background: Dynamical Decoupling
DD can in general suppress generic interactions through pulses that rotate around multiple axis [31]. However, DD is most resource efficient when tailored to the specific type of interactions at hand [32, 33]. Furthermore, the most effective DD sequence depends on the noise type present in the physical system.
@@
$\text{CPMG}\equiv\tau/2-X-\tau-X-\tau/2,$ (1)
where $\tau$ is the duration of the free evolution. Multiple CPMG sequences can be concatenated one after another, see Fig. 2(a). While the CPMG sequence can suppress homogeneous dephasing along one axis, it cannot suppress noise stemming from generic system-environment interactions. By contrast, the XY4 DD sequence [37], defined by
$\text{XY}4\equiv Y-\tau-X-\tau-Y-\tau-X-\tau,$ (2)
and shown in Fig. 2(b), is a universal DD sequence that can suppress generic system-environment interactions. It employs $\pi$-rotations around the $x$ and $y$-axis described by Pauli operators $X$ and $Y$, respectively. However, the effectiveness of these DD sequences critically depends on the noise, i.e., the detrimental interactions that are present in the system, which is often challenging to infer in superconducting quantum devices [23]. As such, the performance of a DD sequence can vary substantially across different device architectures.
Figure 2: Schematic diagram of DD sequences. (a) Two concatenated CPMG sequences. (b) Two concatenated XY4 sequences. (c) the LDD sequence we introduce in this work. All sequences are depicted for an idle period of $4\tau$. CPMG uses only $X$ gates (blue rectangle) whereas XY4 uses both $X$ and $Y$ gates (green rectangles). LDD uses parameterized gates, $R_{\vec{x}^{*}}$ (purple rectangles) where the optimal rotation angles $\vec{x}^{*}$ are found by classical optimization routines performed in tandem with a quantum device. A single LDD sequence is defined by $\tau-R_{\vec{x}^{*}}-\tau-R^{\dagger}_{\vec{x}^{*}}-\tau$, where $\tau$ is the duration of the free evolution.
@@ III Learning how to suppress noise
DD gates themselves are prone to errors, such as errors in the rotation angles. Furthermore, the optimal rotation axis of the gates in the DD sequence may depend on the type of noise in the system. We address these issues by adopting a different optimization approach from Tong et al. [30]. Instead of fixing a set of DD gates and optimizing over the DD sequence structure to find the best DD sequence [30], we optimize the rotational parameters entering in the chosen DD sequence to improve performance. This approach is similar to finding noise resilient quantum circuits through machine learning [39]. After introducing the theory behind such a learning dynamical decoupling approach, we demonstrate on IBM Quantum hardware in two experiments that the optimized LDD sequences yield the best performance (compared to XY4 and CPMG) in suppressing errors.
@@ III.1 Theory
to an initial state $\rho_{0}=|\psi_{0}\rangle\langle\psi_{0}|$, which we assume is pure, to create a desired target state $\rho_{\text{ideal}}=|\psi_{\text{ideal}}\rangle\langle\psi_{\text{ideal}}|$. Here, we concern ourselves with noise described by a collection of typically unknown quantum channels $\mathcal{M}_{j}$, in which we include potential mid-circuit measurements, that act between the unitary channels $\mathcal{U}_{j}$ when qubits are idling. The noisy quantum circuit then takes the form $\prod_{j}\mathcal{M}_{j}\mathcal{U}_{j}$. Each noise channel $\mathcal{M}_{j}$ can describe unitary and non-unitary errors. To suppress noise induced by $\mathcal{M}_{j}$ we divide $\mathcal{M}_{j}$ into $N+1$ channels $\mathcal{M}_{j,k}$ so that $\prod_{k=1}^{N+1}\mathcal{M}_{j,k}=\mathcal{M}_{j}$, and insert a LDD sequence of the form shown in Fig. 2(c). The LDD sequence is described by parameterized unitary quantum channels $\mathcal{R}_{\vec{x}}(\cdot)=R_{\vec{x}}(\cdot)R_{\vec{x}}^{\dagger}$ that are applied between the $\mathcal{M}_{k}$’s where the LDD gates $R_{\vec{x}}$ are parameterized by $\vec{x}\in\mathbb{R}^{M}$. For simplicity, we assume that the LDD parameters $\vec{x}$ are the same for each LDD gate $R_{\vec{x}}$. If we include $N/2$ LDD gates between the noise channels $M_{j,k}$, followed by implementing $N/2$ times the corresponding inverse $R_{\vec{x}}^{\dagger}$ to insure that the total LDD sequence composes to the identity in the absence of noise, as depicted in Fig. 2, the noise channel $\mathcal{M}_{j}$ becomes
$\mathcal{M}_{j,\vec{x}}:=\mathcal{M}_{j,N+1}\prod_{k=N/2}^{N}\mathcal{R}_{\vec{x}}^{\dagger}\mathcal{M}_{j,k}\prod_{k=1}^{N/2}\mathcal{R}_{\vec{x}}\mathcal{M}_{j,k}.$ (4)
@@
To optimize the parameters $\vec{x}$ we need a cost function $J(\vec{x})$ that is sensitive to the quality of the circuits in Eq. (3). For small circuits, a natural choice for a figure of merit or cost function is the fidelity error,
$J(\vec{x})=1-\langle\psi_{\text{ideal}}|\rho(\vec{x})|\psi_{\text{ideal}}\rangle,$ (6)
where $F=\langle\psi_{\text{ideal}}|\rho(\vec{x})|\psi_{\text{ideal}}\rangle$ is the fidelity with respect to the ideal state. However, state tomography scales exponentially with system size. For large circuits, a scalable cost function can be built in multiple ways. As done in Refs. [40, 30] we can invert the circuit with mirroring in which each $\smash{\mathcal{U}_{j}^{\dagger}}$ is applied in reverse such that $\rho_{\text{ideal}}=\rho_{0}$. Alternatively, one can reduce the original quantum circuit to a Clifford circuit as done, for instance, in Refs. [41, 42]. Here, the single-qubit gates in a circuit (with the CNOT as the two-qubit gate) are replaced by Clifford gates such that the whole circuit is a Clifford gate. In hardware where parameters are encoded in virtual-$Z$ gates, this will preserve the structure and timing of the underlying pulses, thereby leaving most noise sources such as $T_{1}$, $T_{2}$ and cross-talk unchanged at the pulse-level. We can then compute $\rho_{\text{ideal}}$ with efficient Clifford based simulators.
@@ III.2 Case studies on IBM hardware
We study the performance of LDD in two different experiments carried out on IBM Quantum hardware with Bell pairs; a valuable resource. For instance, Bell pairs enable quantum gate teleportation [44] and similar known states generated in a factory of resources to enable circuit cutting [22]. In the first experiment we suppress noise during mid-circuit measurements. In the second experiment we suppress noise resulting from an increasing circuit depth. In both cases we minimize fidelity loss due to noise by increasing the fidelity of preparing the Bell state $|\psi_{\text{ideal}}\rangle=|\Phi_{+}\rangle=\frac{1}{\sqrt{2}}(|00\rangle+|11\rangle)$, on two qubits $q_{i}$ and $q_{j}$ within a $n>2$ qubit system, starting from $\smash{\rho_{0}=|0\rangle\langle 0|^{\otimes n}}$. To infer the value of $J(\vec{x})$ in Eq. (6) for the Bell state, it suffices to measure the expectation values of the $X_{i}X_{j}$, $Y_{i}Y_{j}$ and $Z_{i}Z_{j}$ Pauli operators with respect to $\rho(\vec{x})$, rather than performing full state tomography [45, 46, 47, 48, 49, 50, 51], i.e,
@@
We employ parameterized decoupling operations given by arbitrary single-qubit rotations
@@
Below we compare LDD with CPMG and XY4 where DD sequences are inserted when qubits are idling. Throughout the two experiments we concatenate two CPMG and XY4 sequences and adjust the duration $\tau$ to match the idling times of the qubits.

### v2

@@ Abstract
Current quantum computers suffer from noise that stems from interactions between the quantum system that constitutes the quantum device and its environment. These interactions can be suppressed through dynamical decoupling to reduce computational errors. However, the performance of dynamical decoupling depends on the type of the system-environment interactions that are present, which often lack an accurate model in quantum devices. We show that the performance of dynamical decoupling can be improved by optimizing its rotational gates to tailor them to the quantum hardware. We find that compared to canonical decoupling sequences, such as CPMG, XY4, and UR6, the optimized dynamical decoupling sequences yield the best performance in suppressing noise in superconducting qubits. Our work thus enhances existing error suppression methods which helps increase circuit depth and result quality on noisy hardware.
@@ I Introduction
Figure 1: Illustration of LDD. The left and the right block indicate a quantum device and a classical computer, respectively, working in tandem. A DD sequence parameterized by $\vec{x}$ is added to idling qubits in a given quantum circuit, here exemplified by a SWAP gate on $(q_{1},q_{2})$ followed by a CNOT on $(q_{0},q_{1})$. The arbitrary single-qubit rotations $R_{\vec{x}}$ in the DD sequence are realized by three parameterized virtual $Z$-gates (circular arrows) and two $\sqrt{X}$ gates (pulse with in-phase and quadrature shown in light and dark blue, respectively). The noisy output state $\rho(\vec{x})$ is compared to the expected ideal state $\rho_{\text{ideal}}$ through the cost function $J(\vec{x})$. A classical optimization algorithm minimizes $J(\vec{x})$ to find the parameters of the gates in the DD sequence that best fit the circuit.
@@
In recent years, DD has become a major error suppression method for noisy superconducting quantum computers [16]. Indeed, the method is easy to apply as a simple transpilation pass that inserts delays and pulses into a quantum circuit. Furthermore, simple sequences such as $X-X$ already yield excellent results [20]. More elaborate sequences, such as staggered $X-X$ [21] and staggered $XY4$, improve, for instance, the execution of dynamic circuits by cancelling cross-talk [22]. Here $X$ and $Y$ are $\pi$-rotations around the qubit’s $x$ and $y$-axis, respectively. Crucially, the performance of a DD sequence depends on the interactions present in the quantum hardware. In superconducting qubits [23], a good model of these interactions is typically not known, a case familiar to optimal control [24, 25, 26, 27] that can be overcome with closed-loop optimization [28, 29]. Similarly, it is possible to tailor DD sequences to the quantum hardware at hand by learning them through genetic algorithms [30]. However, in hardware, DD pulses are not perfect, potentially introducing additional noise and errors through the controls that implement the DD pulses, thereby diminishing the quality of the designed DD sequence. While control noise-robust DD sequences exist [13], e.g., suppressing errors in rotational angles of the DD gates, such sequences typically require knowledge of the type of control error that is introduced.
@@
To overcome these limitations, in this work we tailor the DD sequences to the hardware and quantum circuits to execute. The goal is to design DD sequences that suppress potentially unknown pulse imperfections and noise. This is achieved by optimizing the rotational angles of the gates in the DD sequence in a closed-loop with the quantum hardware. A classical optimizer is fed the cost function value that is reconstructed from quantum samples and that is sensitive to the quality of the DD pulses, see Fig. 1.
@@
This manuscript is structured as follows. We introduce in Sec. II hree commonly employed DD sequences CPMG, XY4 and UR6. Next, in Sec. III we develop the theoretical framework of how optimal parameters in DD sequences are found on quantum hardware, which we refer to as learning dynamical decoupling (LDD). We demonstrate in Sec. III.2 the utility of LDD on IBM Quantum hardware by comparing the performance of LDD to CPMG, XY4, and UR6 to suppress noise in two experiments. We show that LDD outperforms CPMG, XY4, and UR6 for suppressing noise present during mid-circuit measurements and noise resulting from increasing the depth of a quantum circuit. We conclude in Sec. IV.
@@ II Background: Dynamical Decoupling
DD is a well known strategy that reduces noise by suppressing unwanted interactions with the environment or undesired couplings of the controls with the system. For an overview of DD strategies we refer to [13]. DD can in general suppress generic interactions through pulses that rotate around multiple axis [31]. However, DD is most resource efficient when tailored to the specific type of interactions at hand [32, 33]. Furthermore, the most effective DD sequence depends on the noise type present in the physical system.
Figure 2: Schematic diagram of DD sequences. (a) Two concatenated CPMG sequences. (b) Two concatenated XY4 sequences. (c) One UR6 sequence. (d) Two concatenated LDD sequences. All sequences are depicted for an idle period of $2\tau$ while the gate implementation time is not shown. When running on hardware, the idle times are reduced to account for the duration of the physical pulses of the DD gate. CPMG uses only $X$ gates (blue rectangle) whereas XY4 uses both $X$ and $Y$ gates (green rectangles). Here $X$ and $Y$ are $\pi$-rotations around the qubits $x$-axis and $y$-axis, respectively. UR6 uses $X^{{}^{\prime}}$ gates that consists of two phase gates with angles $32\pi/$ 3 (shown as blue pulses) with opposite rotational directions that are inserted before and after an $X$ gate, see Eq. (3). LDD uses parameterized gates, $R_{\vec{x}^{*}}$ (purple rectangles) where the optimal rotation angles $\vec{x}^{*}$ are found by classical optimization routines performed in tandem with a quantum device.
@@
$\text{CPMG}\equiv\tau/2-X-\tau-X-\tau/2.$ (1)
Here, the total duration of the free evolution is $2\tau$ and $-\tau-$ indicates an idle period of duration $\tau$. Multiple CPMG sequences can be concatenated one after another, see Fig. 2(a). While the CPMG sequence can suppress homogeneous dephasing along one axis, it cannot suppress noise stemming from generic system-environment interactions. By contrast, the XY4 DD sequence [37], defined by
$\text{XY}4\equiv Y-\tau/2-X-\tau/2-Y-\tau/2-X-\tau/2,$ (2)
is a universal DD sequence that can suppress generic system-environment interactions. Two concatenated XY4 sequences are shown in Fig. 2(b). Both CPMG and XY4 are sensitive to pulse imperfections [38]. Such errors can be suppressed with a universal rephasing (UR) DD sequence by inserting properly tailored phase gates into, e.g., a CPMG to obtain the symmetric UR6 DD sequence [39]
$\begin{split}\text{UR}6\equiv{}&2\tau/7-X-2\tau/7-X^{{}^{\prime}}-2\tau/7-X-2\tau/7\\ &-X-2\tau/7-X^{{}^{\prime}}-2\tau/7-X-2\tau/7.\\ \end{split}$ (3)
The UR6 sequence has four $X$ gates and two phased $X$ gates denoted by $X^{{}^{\prime}}=R_{z}\left(-\frac{2\pi}{3}\right)XR_{z}\left(\frac{2\pi}{3}\right)$ where $R_{z}(\theta)$ is a rotation around the qubits $z$-axis by an angle $\theta$.
DD sequences with higher order protection can be built with additional single-qubit pulses, for example, by concatenating existing sequences [40] or considering semi-classical noise models [39]. However, the effectiveness of DD sequences critically depends on the type of noise, i.e., resulting from the detrimental interactions that are present in the system, which is often challenging to infer in superconducting quantum devices [23]. For example, Ref. [12] shows that the performance of different DD sequences is device dependent even though the devices all leverage the cross-resonance interaction [41]. Furthermore, pulse imperfections, such as a detuning in frequency or amplitude, degrade the performance of a DD sequence [39]. As such, the performance of a DD sequence can vary substantially across different devices and even time, if hardware parameters drift.
@@ III Learning how to suppress noise
DD gates themselves are prone to errors, such as errors in the rotation angles. Furthermore, the optimal rotation axis of the gates in the DD sequence may depend on the type of noise in the system. We address these issues by adopting a different optimization approach from Tong et al. [30]. Instead of fixing a set of DD gates and optimizing over the DD sequence structure to find the best DD sequence [30], we optimize the rotational parameters entering in the chosen DD sequence to improve performance. This approach is similar to finding noise resilient quantum circuits through machine learning [43]. After introducing the theory behind such a learning approach to dynamical decoupling, we demonstrate on IBM Quantum hardware in two experiments that the optimized LDD sequences are better at suppressing errors than the CPMG, XY4 and UR6 sequences.
@@ III.1 Theory
to an initial state $\rho_{0}=|\psi_{0}\rangle\langle\psi_{0}|$, which we assume is pure, to create a desired target state $\rho_{\text{ideal}}=|\psi_{\text{ideal}}\rangle\langle\psi_{\text{ideal}}|$. The quantum channels $\mathcal{U}_{j}$ may correspond to the ideal gates in a circuit. Here, we concern ourselves with noise described by a collection of typically unknown quantum channels $\mathcal{M}_{j}$, in which we include potential mid-circuit measurements, that act between the unitary channels $\mathcal{U}_{j}$ when qubits are idling. The noisy quantum circuit then takes the form $\prod_{j}\mathcal{M}_{j}\mathcal{U}_{j}$. Each noise channel $\mathcal{M}_{j}$ can describe unitary and non-unitary errors. To suppress noise induced by $\mathcal{M}_{j}$ we divide $\mathcal{M}_{j}$ into $N+1$ channels $\mathcal{M}_{j,k}$ so that $\prod_{k=1}^{N+1}\mathcal{M}_{j,k}=\mathcal{M}_{j}$, and insert a LDD sequence of the form shown in Fig. 2(c). The LDD sequence is described by parameterized unitary quantum channels $\mathcal{R}_{\vec{x}}(\cdot)=R_{\vec{x}}(\cdot)R_{\vec{x}}^{\dagger}$ that are applied between the $\mathcal{M}_{k}$’s where the LDD gates $R_{\vec{x}}$ are parameterized by $\vec{x}\in\mathbb{R}^{M}$. For simplicity, we assume that the LDD parameters $\vec{x}$ are the same for each LDD gate $R_{\vec{x}}$. If we include $N/2$ LDD gates between the noise channels $M_{j,k}$, followed by implementing $N/2$ times the corresponding inverse $R_{\vec{x}}^{\dagger}$ to insure that the total LDD sequence composes to the identity in the absence of noise, as depicted in Fig. 2, the noise channel $\mathcal{M}_{j}$ becomes
$\mathcal{M}_{j,\vec{x}}:=\mathcal{M}_{j,N+1}\prod_{k=\frac{N}{2}+1}^{N}\mathcal{R}_{\vec{x}}^{\dagger}\mathcal{M}_{j,k}\prod_{k=1}^{N/2}\mathcal{R}_{\vec{x}}\mathcal{M}_{j,k}.$ (5)
@@
To optimize the parameters $\vec{x}$ we need a cost function $J(\vec{x})$ that is sensitive to the quality of the circuits in Eq. (4). For small circuits, a natural choice for a figure of merit or cost function is the fidelity error
$J(\vec{x})=1-\langle\psi_{\text{ideal}}|\rho(\vec{x})|\psi_{\text{ideal}}\rangle.$ (7)
Here, $F=\langle\psi_{\text{ideal}}|\rho(\vec{x})|\psi_{\text{ideal}}\rangle$ is the fidelity with respect to the ideal state. However, state tomography scales exponentially with system size. For large circuits, a scalable cost function can be built in multiple ways. As done in Refs. [44, 30] we can invert the circuit with mirroring in which each $\smash{\mathcal{U}_{j}^{\dagger}}$ is applied in reverse such that $\rho_{\text{ideal}}=\rho_{0}$. Alternatively, one can reduce the original quantum circuit to a Clifford circuit as done, for instance, in Refs. [45, 46]. Here, the single-qubit gates in a circuit (with the CNOT as the two-qubit gate) are replaced by Clifford gates such that the whole circuit is a Clifford gate. In hardware where parameters are encoded in virtual-$Z$ gates [47], this will preserve the structure and timing of the underlying pulses, thereby leaving most noise sources such as $T_{1}$, $T_{2}$ and cross-talk unchanged at the pulse-level. We can then compute $\rho_{\text{ideal}}$ with efficient Clifford based simulators.
@@ III.2 Case studies on IBM hardware
We study the performance of LDD in two different experiments carried out on IBM Quantum hardware with Bell pairs; a valuable resource. For instance, Bell pairs enable quantum gate teleportation [49] and similar known states generated in a factory of resources enable circuit cutting [22]. In the first experiment we suppress noise during mid-circuit measurements. In the second experiment we suppress noise resulting from an increasing circuit depth. In both cases we minimize fidelity loss due to noise by increasing the fidelity of preparing the Bell state $|\psi_{\text{ideal}}\rangle=|\Phi_{+}\rangle=\frac{1}{\sqrt{2}}(|00\rangle+|11\rangle)$, on two qubits $q_{i}$ and $q_{j}$ within a $n>2$ qubit system, starting from $\smash{\rho_{0}=|0\rangle\langle 0|^{\otimes n}}$. To infer the value of $J(\vec{x})$ in Eq. (7) for the Bell state, it suffices to measure the expectation values of the $X_{i}X_{j}$, $Y_{i}Y_{j}$ and $Z_{i}Z_{j}$ Pauli operators with respect to $\rho(\vec{x})$, rather than performing full state tomography [50, 51, 52, 53, 54, 55, 56]. The cost function is thus
@@
The three correlators are each estimated with a quantum circuit executed with 400 shots in each evaluation. We employ parameterized decoupling operations given by arbitrary single-qubit rotations
@@
Below we compare LDD with CPMG, XY4, and UR6 where DD sequences are inserted when qubits are idling. Throughout the two experiments we concatenate two LDD, CPMG and XY4 sequences, as shown in Fig. 2, and adjust the duration $\tau$ to match the idling times of the qubits. Therefore, the LDD sequence is built from two $R_{\vec{x}}$ gates and two $R^{\dagger}_{\vec{x}}$ gates. At the hardware level, each $R_{\vec{x}}$ gate is implemented by three virtual-Z rotations and two $\sqrt{X}$ gates. The $X$ and $Y$ gates are implemented by a single pulse with the same duration as a $\sqrt{X}$ gate. An example of the transpiled DD sequences is shown in Appendix D.